Tet's Web Indexing Thumbnailer
Parent Dir
20040801Arri and the microwave (apparatus for filming the nuking of CDs), holography apparatus (stone table), hologram being reconstructed.
20040816Dimond plate view camera, holography apparatus.
20040919Holography apparatus, VRP-M film processed with GP-2, beam splitter, apparatus to measure diffraction efficiency.
20041011French drain at pimmit house, the imperial net-star at work, arrivial of fume hood, panorama of basement.
20041114dimond plate view camera, film holder.
20041120Pulsed yag laser, disassembly and inspection, playing with trying to get frequency doubler better aligned.
20041122Pulsed yag laser frequency doubler, condensate pump, pulsed laser power supply.
20041207Pulsed laser power supply, holography apparatus, copy arrangement.
20041219Holography apparatus, shots of basement, attempting to take pictures of a hologram.
20041220More attempts to take pictures of a hologram (23.2).
20041230Fabrication of holography plate holder, stuffing argon laser in its bucket, holography apparatus.
20050206Holography apparatus drying after being painted, fabrication of spatial filter,
2005020602Comparision between original spatial filter, and new (functional) version.
20050227Pictures of inside of crawl space at pimmit hills house.
20050404Valve tube extensions on innertubes, for holography apparatus. Fabrication of multiplex film holder.
20050406Continuation of fabrication of multiplex hologram plate holder.
20050509Holography apparatus, sump pump drain at Pimmit Hills, pulsed laser interferometer (with bite marks on beam splitter).
20050529Building the video switch used on my web camera.
20050625Multiplex hologram apparatus, arrival of 3x4 foot breadboard, making vertical threaded rod/flange assemblies, second big lens mount, beam splitter assembly.
20050725Laser stand, cinder block holography table, two beam transmission holography apparatus.
20050821Martin, first version of the can crusher (attempted triggered spark gap).
2005082102First crushed can, badly burned spark gap electrodes (use brass!).
20050828panorama of basement, spark gap with brass electrodes, me in Dr. Horrible goggles (I had them first!), attempt to power 6 foot HeNe laser, MFM hologram.
20050830Thyratron hologram master setup, kludged rainbow hologram setup.
20051010Multiplex hologram plate holder, Arrival of big orange YAG laser head, multiplex hologram camera bracket, camera geometry observations.
20051016More of multiplex hologram camera bracket, test shots of plant.
2005101602Multiplex hologram dataset. 53 pictures of a poinsetta.
20051022Multiplex hologram setup, computers on rack (and my now broken lava lite...).
20051025Reconstructing the multiplex hologram master, with a red laser. Very Dim.
20051202Adding a regulator to the 24V power supply, so its actually 24V, disassembly of rod/lamp assembly from orange YAG laser, box of PFG-3c plates.
20051216Christmas hologram setup, fabrication of magazine adaptor for Tim's film printer.
20051221Voltage regulator board for 24V power supply, Arri stuffed in my van to shoot driving to work on 495 at 6 frames/sec.
20060110Making 1.5 inch cylinder and flange assembly for optical table, 1.5 inch square bar and flange with a zillion hand-threaded holes.
20060117Milling of the big lens mount and beam splitter brackets.
20060212fabrication of beam splitter brackets, big lens mount, H1H2 copy setup, thyratron hologram
20060222painting fort (cardboard, fan with filter), pagent projector innards, bleaching holograms
20060302Experiment re-bleaching holograms and then trying to reduce printout with acetic acid, thyratron, charging power supply,
20060403Mastering hologram of beany lion and xmas ornament (cinderblock optical table), removing oxide from geiger counter, arrival of 16mm printer and light controller.
20060510First version of the Peril variac assembly, test run of the still.
20060617can crusher, assembly of spark gap and dump switch, mysterious card cages from some video processor.
20060618can crusher experiments, voltage ramp on spark gap, crushed rootbeer can.
20060711Rebuilding a pedal splicer, arrivial of the optical table, thyratron test assembly, can crusher charging power supply, innertubes which supported cinderblock optical table.
20060729Charging supply for can crusher, physics cookout, installation of optical table (at Pimmit house).
20060813Physics cookout, fun with the can crusher and cinematography, destroyed diodes and coil.
20061007Edching Dan's LED driver, multiplex plate holder, LED button assembly (first prototype), doomed plexiglass plate holder (for processing, had chemical compatability issues, disintigrated).
20061009stainless steel plate holder (for processing), first successful PFG-3c holograms.
20061112Construction of the Peril variac assembly, super-safelight warming up, replaced tiles on pimmit house, enlarger pieces.
20061126fabrication of second beam splitter.
20061204Three beam mastering setup, focus depth test, second beam splitter
20061210Dan't hologram, Tim's hologram, pouring old JD2 bleach (dichromate) into dead D76R, focus depth test.
20070113Adapting lens onto D-II enlarger, power supply isolation switch for can crusher, pressure regulator, simmer supply experiment.
20070123Kody's mushroom, simmer power supply
20070219Patching air intake filter pot on nissan van, 6V "lantern battery" containing D batteries, assembling still.
20070305Getting the pulse laser, optics with most of protectiving tubing still in place, interior of heat exchanger, weird glass boxen.
20070312Pulse laser power connections, interior of control box, capacitor bank, charging supply, rod/lamp assembly.
20070317Replacing all the O-rings in pulse laser, Dan holding LED staff for Kody.
20070412replacing pulse laser coolant tubing, still, alignment laser, baffles for making test strip holograms
20070426drying negatives, broken glass tubing (thanks UPS), ten oaks house (outside only), other house on ten oaks, really impressive CCD scan artifact.
20070508damage to van from drunk driver, house hunting: hilltop house.
20070510House hunting: ten oaks house, others.
20070511House hunting: really scary house in Savage, hilltop house, others.
20070517Pulse laser innards, living room (Pimmit house), Dan sorting parts, playing with Dan's coherent 315m and print washer, traffic.
20070524Shiny truck, motor cycles, attempts to get CCD scan artifact.
20070604Showing how to load federal 4x5 processing tank, traffic, yard of ten oaks house.
20070610traffic accident, flooding on 495, repair of capacitor bank, cement things attached to tree in yard.
20070620Traffic on 495, traffic in rain, test-fit of big screen in living room, van in garage, putting lid on septic tank, shots of yard/cemetary.
20070704Moving pulse laser to ten oaks, fios install, microscopes, traffic accident
20070729Shelves at ten-oaks, traffic accident, moving furnature, laser in basement, drive to Iowa
20070805Carpets in van, 1 light second, moving to ten oaks, moving the optical table
20070812Reel to reel, John and heavy guitar, basement full of stuff, carpet, stuff filling all the rooms, moving to ten oaks.
20070823Basement of Ten Oaks house, weird old switches, night pictures on 29
20070826Back yard, green on side of house, lots of pictures of Pimmit house (empty)
20070906Pictures on way to MFM, stuff in garage, pressure-washing basement floor.
20070919Failed epoxy paint for basement floor, successful epoxy paint, epoxied sink, interesting car on 29, bunch of pictures of Pimmit house, new boiler, new AC ducts.
20070921Ducts in film room closet, air handler in empty basement, optical table legs, stuff piled in basement.
20070923Stuff piled in basement, deployment of optical table, ducts in knee wall space, dining room, cleaning out well-head, shop-vac.
20071007Epoxy painting floor in basement, moving captive air tank, Dan in his responsible grownup costume, mouse in trashcan, painting basement walls.
20071016Epoxy-painting the basement floor, new boiler, moved captive air tank and filter, optical table in final location, junque in garage.
20071022First batch o' photo chemistry in the new place, installing basement lights, nikor 4x5 tank.
20071105Sorting cables in living room, stuffing van in garage, darkroom curtains, driving to NJ to ge the darkroom sink, assembling pneumatics for optical table.
20071113Holography apparatus (some assembly required), darkroom sink installation, humidifier installation.
20071125Humidifier installation, darkroom light-proofing, holography apparatus, Mom's attic, removal of the old galvanized drain pipe.
20071202Plumbing for the kitchen sink , darkroom sink, and cement sink, holography apparatus, long exposures on 29, giant coke cap project.
20071208Projector cart, Assembly of black slab legs, new red shelf, prints drying, darkroom apparatus.
20071229Birds in Tyson's corner mall, living room, still, bathroom floor vent work.
20080124The still running, new window blinds, weird bug, Xmas cactus blooming, pie, holography apparatus, floor vent in bathroom, lighting controler innards.
20080207Toilet mantainance, transmission holograms, dead bird in basement, pulse laser alignment apparatus, frog in basement window.
20080209Pulse laser firing, zappit paper.
20080303Pulse laser PCboards, old van, pulse laser, pulse laser firing viewed with IR viewer,
20080310Water following the cracks in the garage floor, traffic on 32, rainbow in inner harbor, KJ and grill, pulse laser, shiny new van.
20080317Passive qswitch for pulse laser, apparatus for determining fast axis of wave plates, pulse laser firing, silicone around power line,
20080327Holography apparatus, for Russ's birthday hologram.
20080401Pictures of Russ's birthday hologram, still boiler glowing, fun with ethernet wiring, shiny new osciloscope.
20080407Fun with florescense and florascin.
20080430Pulse laser and osciloscope, gutter, shiney new lawn mower, fun with ducts, pulse laser interferometer and holography apparatus, doodah parade.
20080506Jacob's ladder, camping at Sporty's (beware the durian).
20080527Merangue and lasers, fabrication and installation of my microwave cart, prints drying.
20080602Tigers at the national zoo.
20080616Garage, after pressure washing and then painting with drylock, yard in rain, making of yard stick hologram.
20080627Yard stick hologram, freshly painted porch, arrival of quantel laser parts.
20080731Innards of Quantel capacitor bank, charging power supply, cross section of lamp power cable, high reflector and aperature of pulse laser, moved garage floresent lamp, pulse laser aperature, holography apparatus, slide for Mom.
20080810A slide for Mom, light trap in AC duct, gel on hotplate, meyer bar, Kody and drill press, can crusher version 3, aftermath of coin crushing coil exploding.
20080827Several failed dichromated gel holograms, yardstick hologram, spraypainting
20080907Lots of attempts to get a good shot of a hologram, apparatus used to take pictures of the holograms, implements of B&W processing on washing machine.
2008090701Lots of attempts to get a good picture of holograms. Artificial light.
20080913Lots of repeated pictures of holograms, in sunlight on the floor.
20081029Failed dichromated gel holograms (note crystals on surface), coin crushing coil assembly, meyer bar, repackaged LCD projector guts, humidifier drain.
20081117Horde of flash bulbs, holography apparatus, toothy critter I got for Xmas.
20090118Test shots for field of view, telecine project (modded pagent), hacdc basement room progress, including abandoned boilers.
20090201Telecine project (modded pagent projector), Dan's LED of Doom.
20090217Pagent shutter, HACDC basement room progress, xenon short arc lamp
20090227My keyboard, painting in the hacdc basement room.
20090309hacdc basement space painting the walls, phone lines on exterior walls, living room in sanctuary, Plumbing of Doom.
20090314hacdc basement space etching the floor, church roof.
20090315Very blury (as opposed to fuzzy) squirells, playing with new projector
20090322Floor paint in hacdc's basement room, moving the plexiglass.
20090325SP-lamp-lp1 xenon lamp for an infocus 130 DLP projector.
20090328Timonium hamfest, playing with lasers in the fog at night.
20090404super speed graphic bits, moving the 6 inch plexiglass part 2 (completion)
20090420Accident on 29, can crusher relays, Doodah parade, Ocean City NJ, accident on 95.
200905022009 physics cookout.
20090510DC traffic accident, 2008 invention of the year cert, building hacdc Shelf of Doom (tm).
20090531Pulse laser, attempting to install second amplifier, building projector stand, crushing coins, mangled solenoid.
20090614Projector on stand, computer mantainance, big turtle, broken tree branches.
20090620Finaly got the IR interface working on my LED button, and cleaned up the branch which fell last saturday.
20090627Arizona/California, yet another broken branch in Dayton.
20090712All these files are corrupt, thanks to the wonders of USB.
20090726Fabrication of Dan's steam punk monitor, laser coax power cable, and a stained glass crab at BWI.
20090802Pictures of quantel's manufactured connector, fabrication of my replacement.
20090808Adding second amp to pulse laser, 220V wiring and drilling/tapping holes in optical rail.
20090823Hanging up my antenna cable, filming my pulse laser firing.
20090919Aero Ektar lens, open-moko surgery, MFM-13, ren faire.
20091006Dan's diode and YAG assembly, well maintanance, Telecine project test data.
20091025Web camera, telecine project.
20091030Telecine project, Orthicon tube, 1886 issue of Scientific American Supplement.
20091106BigC fix on my openmoko, Telecine project.
20091107Flashbulbs, something Z80 based.
20091209FoxLedge brand distilled water, mown leaves, insulated hydronic lines, exploded LCD monitor, HacDC's reprap machine.
20091218exploded LCD display, HacDC's reprap, freshly fabricated saw horses, construction of Mom's TV stand.
20100109Lots of snow, LED button soldering, fabrication of telecine main board.
20100118Telecine project: LED lamphouse, encoding wheel.
20100129New Teryx, telecine project, Dan's light (Of Doom).
20100206Telecine, featuring a Nikon D300 (they make good glass, and lousy firmware), and some shots of the snowpocolypse.
20100207Snowpocalypse: the excavation.
20100212Telecine motor pulley, snow and icecicles.
20100214Terrier missile antenna, swap-fest at national electronics museum.
20100306Holography, gnomon, aero ektar and the UV thing, $13 wrench.
20100309Reprap machine down at HacDC.
20100319Lots of rain, to observe low spots in yard, gnomon, aero-ektar lens being bleached, holograms.
20100408Reprap machine at HacDC, it finaly finished a gcode file, laser and harddisk at Timonium hamfest, installing trees, aero-ektar lens being bleached cont.
20100419Aero Ektar lens being bleached, 2010 DooDah parade.
20100508Making banana slugs at HacDC, repairing a neon sign, Ash trying to light his cigarete, adding outlets to HacDC's basement room.
20100516The old patent office library, Joint Services Open House.
20100524Messing with a $20 bill, color perception experiment.
201005292010 physics cookout. (no injuries!)
20100601Dan's brown-n-orange, DMD terminal, reshoot of nitrate film.
20100621silver-nitrate stains, darkroom plumbing, frequency doubled YAG through a prism, through an IR viewer (showing 808, 1064, and 532nm), holograms 67.[1-5].
20100702Finally got enough 532nm photons to put a mark on zapit paper. New optics, glassware, film.
20100706Innards of a magnetic stirer, pulsed YAG, metro, HacDC workshop and reprap.
20100718Dry fruit trees, hanging hacdc projector, shark building, 45deg mount for polarizer, survey stakes, titanium, pulse laser tweeking, IR LEDs in oscilloscope.
20100724More shark building, pulsed holography attempts (holograms 68.1 to 68.6).
20100725Pulsed holography: a bunch of Denisyuk attempts. (reconstructed with a laser)
20100808Pulsed holography attempts: 2 beam transmission holograms 69.1 and 69.2. (I'm using the digital camera to look at where the non-eye-safe laser is shining.)
20100828Sorting prints, new washing machine and dryer, quantel laser circuitry, putting brass end on 4 gauge wire, looking for arc in quantel laser.
20100829Fiddling with secondary lens on digital camera, looking for source of weird popping sound. Seems to be an arc to gnd on the second amp.
2010082901Looking for source of weird popping sound. Seems to be an arc to gnd on the second amp.
20100831Fun with a laser pointer with a diffraction grating, and long exposures.
20100907Dave's laser pointer, and IR leak detection apparatus.
20100930HacDC, holography apparatus (new mirror mount, new lens mount), old dryer.
20101003More laser pointers, IR viewers, and prisms. (and calibrated postit)
20101017Avalanche transistors, 473nm laser shutter assembly, Ren Faire.
20101029Reprap, rest of 473nm laser shutter assembly.
20101030Rally to restore sanity and/or fear.
20101123Reprap gets a new bearing, fabrication of qswitch trigger PCB.
20101204Driving home from Blacksburg, more qswitch trigger fabrication.
20101205More qswitch trigger fun. It worked for the first time today.
20101208Reference pictures of SHV, MHV, and BNC connectors.
20101211More pictures of the quantel laser, aligning it with holography apparatus.
20101215Couple of shots of qswitch on quantel, printfinder laser.
20101218Qswitch trigger and power supply, finaly fully assembled.
20101221Final assembly of the qswitch trigger, attempts to take pictures of the lunar eclipse, mysterious thing which is labeled mirror, and little telescope.
20101229Arrivial of the Lumonics ruby, bunch of interior shots.
20101231Bunch of pictures of the Lumonics ruby laser.
20110108Lumonics laser coolant fun, last pictures of P260 monitors.
20110118Pictures of the Lumonics laser after unpleasantness turning it on the first time.
20110126Lumonics laser's alignment laser.
20110206Bunch of Lumonics innards, kludged original controller, first voltage in the OSC capacitor.
20110212Working on Lumonics control panel.
20110219Completed Lumonics control panel. (of course its still empty...)
20110227Replacement of well pump, etching and drilling of Lumonics controller PCB.
20110303Lumonics laser controller test power supply, hologram lamps, more PCB assembly.
20110306More Lumonics controller assembly, panel wiring, first Application of Power.. Girl Power at APL.
20110313Even more lumonics ruby laser assembly. First light.
20110325More lumonics ruby laser assembly, M35 pin holder and soldering, controller board nearly complete.
20110326Lumonics controller fully populated, couple of shots of the amplifier.
20110403More Lumonics ruby laser, measuring the cap voltages, control panel, both osc and amp firing. (and two images lining up the telecine.)
20110409Camera bracket on telecine project, aligning lumonics laser, first coherent light from osc.
20110410Aligning lumonics laser (lots of paper disks) interior of etalon oven.
20110430Adding second coolant loop to lumonics laser, Kes's hologram, Robotfest.
20110508Ice in the lumonics laser's secondary coolant loop, first coherent light through amp.
20110526Telecine power supply, stakes for alleged building, milling on drillpress, DMD running layers.
201105282011 Physics Cookout. (no injuries)
20110530Playing with the can crusher. Used up an aligator clip...
20110605Working on the telecine's takeup mechanism.
20110611Test holograms with geola's sample film.
20110624Telecine, lumonics qswitch trigger board, mode hopping measurement apparatus.
20110627Telecine stepper driver boards, fiberoptics for ruby qswitch trigger, mode hopping measurement, 315M with switching power supply.
20110703315M with switching power supply installed, telecine, 315M with new cover, C532 cavity.
20110709PD2000 board from lumonics, several zillion images of C532 laser, resonator removed from box. First version of adjustment apparatus.
20110724Several zillion more images of C532 laser.
20110829A couple of shots of C532 laser, making of Dave's hologram, damage from the hurricane.
20111002Ren Faire, Solar Decathalon, next version of C532 lens adjuster (3 I think).
20111003Couple of shots of the C532 running.
20111008More shots of the C532, test shots, oil tank valve.
20111022Test shots, Oktoberfest party at Ryan's, electric cars, C532 current measurement apparatus.
20111119Test shots, C532 laser, optical fiber in resonator of ruby laser, qswitch driver board assembly, quantel qswitch driver.
20111124Test shots, lumonics qswitch driver, redoing lumonics power connection, UV LED optical cement curer.
20111125bad pixel shots, avalanche transistors avalanching, cementing output lens back into C532, brackets for indicator stands, badge.
20111203Tent over optical table, C532 laser, new mirror mount.
20111205Pictures of kiln.
20111223532 lasing outside of its case, cool cauliflower, lumonics laser q-switch driver.
20111231Someone pulled over infront of my house, ice crystals, weird spatial filter artifacts.
20120115Enlarger with my great uncle's WW2 patch, oil tank valve replacement.
20120219Plotting PCB artwork (doesn't work, only sharpies survive the etchant) replacing temp sensors, copper based etchant, lumonics laser qswitch driver.
20120301Copper based etchant, lumonics laser qswitch, thermister calibration. (RIP thermometer)
20120305Lumonics laser, optical fibers in rod assemblies, kludged qswitch.
20120318Lumonics laser qswitch driver assembly, Johnny's climbing gym.
20120327Lumonics laser. Attempting to take picture of the maltese cross.
20120408Tree surgury, tire mark from some drunk moron missing the road (we found the bottle too), hologram frame, lumonics laser.
20120415Lumonics laser third alignment laser, head electronics all assembled, and zapit paper taped to the wall. Some holograms which don't suck.
20120429aligning the lumonics laser's spatial filter, Charle's wedding, lumonics laser's lid, Johnny's hologram.
20120505Putting smaller axial mode aperature in the lumonics laser, putting the optical rail into the new enclosure, enlarged spot before amp, external shots of new enclosure.
2012050501lumonics laser spot broadened, before the amp. To determine axial mode structure.
2012050502lumonics laser spot broadened to determine axial mode.
20120513printer parts, branch picnic at APL.
20120519Coin squisher at the 2012 physics cookout (no injuries).
20120519_CFLPictures taken by CFL at the 2012 physics cookout.
20120602Started construction of pulsed apparatus for lumonics ruby laser, bunch of alignment images.
20120603Lumonics laser, pulsed holography apparatus. Attempting to align it.
20120612More lumonics laser, aligning the pulsed holography. First successfull ruby pulsed hologram.
20120615Unhappy hydronic heat line vents.
20120616Lumonics laser, first really successful pulsed hologram, and apparatus.
20120702Damage from the derecho. Start of kitchen work. Everything stuffed into living room, exterior wall interiors.
20120707More kitchen work. completion of exterior work, fixed garage door.
20120712More kitchen work, removal of the old galvanized steel drain pipe. Pictures behind knee wall.
20120715More kitchen work, drywall, sink plumbing in the wrong place, expand-o-foam in the cinderblock wall.
20120729More kitchen work, plumbing for the sink hopfully in the right place this time.
20120805More kitchen work, applying primer. Black snake on the porch.
20120811More kitchen work. Finished painting, unwrapped everything, dining room returns.
20120818Lots of lumonics laser alignment shots, and retroreflection ionizing the air.
20120904Behind Johnny's climbing wall, kitchen in boxen on the porch.
20120909Lamp adaptor in a pageant projector, kitchen work.
20120911More kitchen work.
20120916Crimper's Unhinged.
20120924More kitchen work. Its almost finished...
20120930More kitchen work: first dishes in dishwasher. Angie's Fish.
20121022Ren faire, angies fish, lapping the ways on the mill.
20121109After Sandy (only minor damage, my trees are out of limbs) lapping the ways, and getting gunk out of a ball nut on the mill.
20121120Mill test run, and lumonics heat exchanger fabrication.
20121121Lumonics laser heat exchanger/TEC assembly.
20121122Lumonics laser heat exchanger assembly being tested.
20121208Lumonics laser controler board V2, heat exchanger being integrated into the chiller.
20121223Alignment images for the lumonics laser, and some zappit paper.
20121225Alignment shots for the lumonics laser, cardboard plate holder adaptor, lots of pictures of the resulting holograms.
20130128Lumonics laser pulsed apparatus, adding a green laser for safelight.
20130203Lumonics laser holography, first really good 8x10 master.
201302168x10 image planed hologram apparatus.
20130225Images of first really good 8x10 hologram, Dave's master hologram.
20130302Water baloon test shots, Dave's hologram being copied.
20130304water baloon test shots, fried mirror on lumonics apparatus.
20130306Couple of test shots of water baloon popping, and reconstructed master hologram.
20130310Pictures of finished (H2) holograms of hamster, legos, quartz, crown graphic, Mom, hasselblad.
20130319Girl Power! 2013, replacement mirror in lumonics laser pulsed apparatus, test shots.
20130422Lathe refit, interferometer, tint of Ultimate plates, H2 spoiled by motion.
20130427Alignment shot of Brian, heat exchanger fun, new paper safe repairs, arrivial of argon lasers.
20130502Assembly images of CNC lathe, TD wake.
20130504Kinetic sculpture race, water bomb hologram H2.
20130505Holograms of water baloon, Dave, Brian.
20130512grinding the ball screw at Tim's, Lathe, water baloon hologram.
2013051201Video of hologram of waterbaloon.
20130525Dan plucking his eye out, lumonics laser chiller, Joel's wedding.
20130527Lathe cutting a roller (sortof), geodesic dome.
20130603Jacob's ladder.
201307048x10 hologramy copy setup, lumonics laser alignment tests, master and copy of corn sheller hologram, interior of AO qswitch driver, fruit trees.
20130707More pictures of H2 holograms, parging back of house, garage door brackets, bees.
20130727Dumplings of sauron, lathe Z axis, new C315M laser.
20130806Kody's birthday party.
20130812finished parging back of house, lathe chewing on some aluminum.
20130825More lathe fun, second version of X ball screw mount, shiming tools so the
20130909Lots of bees, motor, poison ivy fruit, lathe
20130917Thirs version of X axis ball screw for the lathe.
20131027Interior of an induction motor (all the better to get it back togather), pulsed holography alignment tests.
20131113Random pictures of my old C315 laser put back togather as a display laser, Dan's rockets.
20131118Alignment tests for holgrams of Dan, my brother, and my mom.
20131202Creepy coolant at Tim's place, long exposures at Charle's parent's place.
20131209Stuffing quantel laser bits in the van, making PCBs with the blue glop.
20131230lumonics laser alignment images, oscilloscope battery, video of squirrel.
20140126Hologram of flower.
20140221New edison motherboard installation, CW yag burning a hole in a 2x4.
20140307Bee hives, Kirk and Ratchel's alignment images, cryocooler.
20140317Cryocooler PCB reference pictures, and qswitched YAG chewing on a 2x4.
20140324More cryocooler pictures, arrival of big dewar.
20140427More cryocooler and dewar integration, bee hives, unhappy scanner, plants in cups.
201405032014 Kinetic Sculpture Race.
20140520Splitting the metaltop hive.
20140525Memorial day coin crushing and holography.
20140608Ryan's birthday party, populating scanner PCB.
20140621plants on porch, daylillies, resolution chart fun, making epoxy coils, 2014 physics cookout.
20140628Assembling the new qswitch trigger for the lumonics laser.
20140705The garden, new qswitch trigger board in the Lumonics laser.
20140713Water melons and corn in the garden.
20140726Water melon not quite ripening.
20140805Holograms of rhododendron, JBadger, Sherbert.
20140810The ground stumps of my dead trees, hula-shower, attempting to observe beam
20140817Attempting the etalon trick on the lumonics laser, evaluating puddle infront of garage.
20140921Yet another attempt at getting the cryocooler to actually accomplish something.
20141006Putting polarizing cube beam splitter on pulsed holography apparatus
201411022014 MD Ren Faire.
20141110Showing movies at Crimper's.
20141202Dramatic sink shot, bee hives, myth busters.
20150103Puting wave plate on Lumonics apparatus, cryocooler attempt, cleaning keyswitch on tandy 200, QR code attempt.
20150126Lumonics apparatus, shmoocon hologram.
20150207Ellicot City train station, Lincoln visit.
20150315Interior of bobdbob, Bee hives (no one home).
20150419Working on the gutter, the hook, replacing ethernet patch cables, day lillies.
20150426Lots of bee hive interiors, cleaning them out for the new bees. Bees in the kitchen.
20150503Extracting honey, cion, intervalometer looking out the window, finishing the paint in the upstairs bathroom.
20150509Pre-physics-cookout equipment shakedown.
20150509_CFLCharles' imagery from the pre-physics-cookout equipment shakedown.
20150516_CFLCharle's imagery from the 2015 physics cookout.
20150525Intervalometer for cookout, mushrooms, final gutter work, filtering honey.
20150601The horde of honey in its jars, and a bunch of pulsed apparatus alignment images.
20150617The convertable we rented in new hampshire, manchester's verizon center, pictures of the six hives.
20150621Assembling the giant pile of hive bodys and supers, installing them, and honey in jar with label.
20150624Headlight bits and skid mark from the asshole who hit the utility pole in my front yard.
20150706Aligning lumonics pulsed apparatus (with milk jug), drain in front of garage, serving size of 50lb bag of sugar.
20150714Lots of random stuff off phone camera.
20151205Redoing the drain in front of the garage, hologram of cat, watermelons, air show, smithsonian with JBadger, hologram of skulls, movies at crimbers, more ram for edison.
20151231scanner, edison ram upgrade, cookies, cat and bee holograms.
20160119Hives after big snow, cryocooler not working, skulls and bees holograms.
20160125Lots of snow, black plastic experiment
20160207broken hand
20160522broken hand, third swarm, toolbox.
20160522_ryans_birthdayRyan's birthday at the outer banks.
20160711Bees, extracting honey, milling machine, Mom's hologram
20160806H2 of Mom's hologram, building mill table, HOPE, moving Ryan.
20160828Assembling the next version of spud, bunch of work on the mill, back of the camera on prototype 3
20161002More assembly of the mill, lifting it onto the table, drains, pumpkin flinger.
20161106Construction of popcorn, and more assembly of the CNC mill.
20161117My milling machine in its box, first flood cooling run, lathe at makertech.
20161127Scanner bits, redone mill drain and Y axis motor shield, scanner computer.
20161128Well repairs.
20161210Lathe manual page with RS-358, fabrication of lamphouse, LED driver chips on LCD.
20161231Popcorn (the computer), baking cookies, bulb for lathe, pillow blocks, scanner.
20170108Control panel, wiring, pillow block ends, lathe, couple of pages from manual, new jaws.
20170210some work on the mill, the scanner, the scanner computer, Lathe lightbulbs, flash gordon.
20170219de-gloping the lathe's coolant tank, film scanner gates, a flat bone for the flat dog at work.
20170305Scanner gates in their foam, grinding the capstan, Mike and Dan at makertech, installation of beaglebone in Ebo's lathe.
20170325More wiring on Ebo's lathe, new coolant pump on my mill, bunch of pictures of the scanner.
20170402Schlieren photography setup, arrival of the bees, pictures of the scanner.
20170403Lots of pictures of the scanner.
201704252017 Physics cookout.
20170521film scanner, 1 shot from FTM, kinetic sculpture race, NS Savannah
20170603film scanner, New mirror mount, bunch of backyard shots, recelling UPS, new bobdbob.
20170627cryogenic coin crushing.
20170812Turning a non-imaging optic, bunch of catalog references, laser cutter innards (and dirty mirrors), aligning ruby laser (put the spatial filter pinhole in).
2017081201Liquid nitrogen, metal detector, cryocooler, filter adapter, ruby laser alignment laser.
20170827Eclipse trek to TN, Chas's sculpture, FTM notes.
20170924Whole bunch of alignment images.
2017092401Cup which Twitchyfox gave me, milling machine making super8 gate, images of H2 of hologram 138.1 being made.
20171021Documentation of moving the spatial filter of the lumonics laser attempting to use more of the amp volume.
20171025Pictures of the tool finder on the lathe, lumonics laser innards.
20171029Pictures of some holograms, test images getting the libusb capture-and-download thing working.
2017102901lumonics end of the osc details, dump relay, dump resistor, alignment and energy measurements.
20171119Whole bunch of pictures of the insides of the (lumonics) JK ruby laser. Boilerplate ruby rods.
20171208boilerplate ruby rod assembly for pump chamber, boilerplate flash lamps and rod in the pump chamber.
20180107More assembly of the pump chamber, coolant connection blocks, assembly of the controller board.
20180114Pump chamber, more assembly, reference images of back of house, pile of laser bits in living room, images from testing flash adapter, laser beam on alignment paper.
2018011401Pulsed apparatus alignment images.
20180122Making of Rain's hologram, laser control parts drying.
20180202Rain's hologram, scale images of LPC4337 dev board, populating laser control PCB.
20180204more assembly pictures of the new laser controller.
20180216Assembling laser control panel, well and sewer cleanout, trigger board assembly.
20180217Lots of attempts to get picture of trigger board test.
20180218Mechanical assembly of new laser control panel.
20180303Wind storm dammage. Need some new trees. :-P
2018030301Makertech move, including how to pick up the lathe.
20180311Some fun behind the knee walls, HOPE holographic apparatus.
2018031101Pictures of holograms, GirlPower! 2018
20180428Building the shutter for the blue laser, removal of my broken spruce tree, ruby laser amplifier power supply construction.
2018042801Picture of mysterious inductor, coin crusher, ruby laser amplifier power supply construction, kaleidoscope.
20180505Dinking around with IR camera.
2018050501Kinetic scuplture race, and one picture of the protractor.
20180506More dinking around with the IR camera, working on the ruby laser amplifier power supply.
20180507Work on the amp power supply, arrival of Square 1.
20180511Images aligning the pockel cell of lumonics laser, showing the maltise cross.
20180519More working on the lumonics laser. Etalon removed, analysing etalon, second amp power supply.
2018051901Thermal pictures of All The Things.
20180526More pictures from the NS Savannah.
20180528Mounting the ruby rod, laser pump chamber assembly, adding third and fourth channels to the qswitch controller.
2018052801More dinking around with thermal camera, warm water running off the driveway.
20180602First light from the first flash lamp in the new amplifier for the lumonics laser.
20180609Lots of pictures of putting the lamps and ruby into the pump chamber. Bunch of pictures of the flashlamp going off. Pictures of the power supply firing, looking for arcing.
2018060901Alignment picture of Red, bunch of pictures of the flash lamp powersupply firing looking for arcing.
20180618Broken inductor, trigger transformer on amp, innards of new teryx, portable holography apparatus, more shots of trigger transformer.
20180624Rehersal for holography workshop, hologram 157.3
20180701Airforce academy, Peterson airforce base, making of hologram 158.[1-3]
20180702Hologram 158.3
20180715exposure measurements on pulsed apparatus, initial test-fit of second amp to optical rail, several test shots of laser and un-connected amp, fully populated power supply (both caps etc.), spatial filter (needs springs).
20180723HOPE 2018 table, and scoping around for possible darkrooms.
2018072301First light through the second amp. The flashlamps went off 7 milliseconds late, so its really not working.
20180726Pictures of the two amp laser control panel, second amp, and new web camera.
20180812Bunch of dinking with IR camera. Test files from correcting BMP file IO.
20180820Bunch of pictures of the ruby laser firing with the second amp. Assembled the power supply onto a dolly, allegedly rated for 1000 pounds and definately not the happy camper.
20180909Hole in the packaging from a poorly packed box from Adorama.
20180930Attempt to take picture of tarantula molt hologram, assembling alumina pump chamber, disassembly and yellowing of alumina, first attempt at combining blue and green lasers.
20181007Lapping the Y axis of my mill.
20181015Lapping the X axis of my mill.
20181104Two pictures taken at the ren faire.
2018110401Fabricating second pump chamber for ruby amplifier, bunch of testing pictures, setup for 2-color holography
2018110402Bee hives, house in the cold, New heatsink arrangement for the ruby laser cooler.
20181118Assembly of peltier heat exchanger, single point GND, bunch of pictures of the beam profile on joulemeter, coolant leak, bunch of pictures of power supply with trigger transformers screwwed down.
20181202Fabricating stainless steel pump chamber endcaps for the ruby laser.
20181208Stainless steel endcaps on pump chamber, bunch of pictures of the amp firing, pondering where the light is escaping, bunch of pictures of the power supply firing, looking for the source of the snapping sound when it fires. Pulsed apparatus, adding side mirror mounts, Optical fiber bracket.
20181211Square One, with christmas lights.
20181222Assembly of the optical rail extension, almost final coolant lines, final power lines, pondering alignment of amp1 with amp0. The beam is hitting the top of something. lumonics
20181231Touristing in DC with Mom and Johnny.
20190101Lumonics case extension, bunch of alignment images.
20190112IR pictures of new chiller heatsinks, boiler, milling machine, exterior of house, Mom's place (test images), new chiller controls, exterior of house (daytime).
2019011201Rebuilding the chiller for the lumonics laser. Its properly thermostaticly controlled now.
2019011202Garden of the gods, museum of mining (lots of steam engines).
20190113Redoing coolant lines on the lumonics laser, holography panel equipment.
20190120CWRU film society marathon number 44.
20190127Installing the colant lines for amp1 on the lumonics laser, then alignment pictures.
20190202Pictures of the defect in the amp1 ruby, alignment pictures of the osc (multiple transverse modes).
20190203Bunch of alignment pictures for the lumonics/JK laser.
20190208Bunch of pictures of holograms, and their subjects.
20190209More alignment pictures of the lumonics laser.
2019020901Alignment of the pulsed apparatus pictures.
20190215Lots of dammaged trees from the ice storm on tuesday or so.
20190217even more alignment pictures from the pulsed apparatus lumonics.
20190220More alignment images. Then the trigger transformer blew up, so pictures of the failed transformer.
20190224repairing the blue laser, pictures of 2 color holograms.
20190310some random lasers, fun with epoxy, and Girl Power!
20190318Installing in the new potted trigger transformers on the ruby laser, bunch of alignment images, beam profile images.
20190323Burned O-rings, new stainless mirror finish reflectors, ruby laser alignment images.
20190324More alignment pictures, hologram 167.1, first viable transmission master on U25.
20190406More fun with the lumonics laser: potting 330 uH trigger transformers, alignment images, striations from the soviet ruby
20190413duplicate image from bifringence in ruby, alignment images
20190414More lumonics ruby laser alignment images.
20190420even more alignment images, lighting tests
20190421Horde of rubies, showing extinction of light in crossed polarizers through ruby, alignment images.
20190428Big Spider, lumonics ruby laser alignment tests and pictures of reconstructed holograms.
2019042801IR images of trigger board.
201905102019 joint services air show
20190512Alignment images for Jim's hologram.
2019051201IR imagery from 2019 physics cookout.
20190519NS Savannah, 2019
20190525Alignment images for Snafu hologram.
20190601Doing H1H2's of Snafu and Jim, then a bunch of RGB laser alignment fun.
20190602A fox in my back yard, then a whole bunch of pictures of H2's.
20190621Arrival of new 830uf caps for the lumonics laser.
20190629Waveforms of capacitor discharge on amp1 of lumonics laser.
20190630Arrival of second lumonics ruby laser, ice in the airconditioner, cooling contraption in sink, color holography object, alignment images of pulsed apparatus and RGB, spectroscopy
20190707Adjusting frame to hold new giant lens, new 532nm safelight, aligning left light source of pulsed apparatus, sketchy spectrograph.
20190721Lots of water in the yard, putting honey in jars.
20190817Bunch of IR pictures taken by Charles.
20190826Dyeing my spotted pants, interior of bobdbob, teryx, spatial filter, rolly thing for laser power supply, putting capacitors on it.
20190827Kinetic sculpture work, ducts at crimber's, projection screen at crimper's.
20190922Pictures of old AC and oil fired boiler.
20190927Climbing with Johnny and Danielle, testing tensionboard.
20191006Drilling appartus at Scott's place, holography panel at FAU.
20191012Drilling the geothermal heatpump well. (made it to 500' there was no water...)
20191013Moving the rack, to get it out of the HVAC guy's way.
20191015Moving the computer, setting up a tent before the driller guys come and drill through the wall.
2019101502Trenching from the geothermal well to the basement.
20191101Installation of geothermal heatpump.
2019110101IR images of geothermal heatpump heating.
20191123AC dampers, heatpump innards, painting floor where oil tank and boiler used to be, stuph jammed into darkroom.
2019112301More painting the floor stuff, fixture for cutting holotimer case, holotimer in its shiny new case.
20191203ph meter, water equipment, test pattern on thermostat.
20191207holotimers, and circuit breakers.
20191222Thermal images of basement, pondering effectiveness of blanket on the wall.
2019122201Pie, and tension board apparatus.
20191228Observing capacitor discharge time of the two caps on the second amp of the lumonics ruby laser after rearranging the electrical connections and bolting everything down.
20191231Aligning the lumonics ruby laser and pulsed holography apparatus from the spatial filter all the way to the plate.
20200124Rules posters for marathon 45.
20200126Diagram of NS Savannah nuclear plant.
20200202aligning the pulsed apparatus, LED cube assembly, first aprilair air filter replacement note UV dammage.
20200223Assembly of LED cube, initial testing.
20200314lumonics pulsed apparatus alignment.
20200406Reference images of back yard for deck project.
20200412Hawk in the tree, plate coating apparatus, exceptionally poorly coated plates,
20200419More dichromated gell holography attempts.
20200426More dichromated gel hologram attempts. Hologram 178.1 to 178.5.
20200427More dicromated gel hologram attempts.
20200429Even more dichromated gel hologram attempts.
20200430More dichromated gel holograms 180.x , broken glass
20200501Even more dichromated gel holograms 181.x
20200502More dichromate holograms, 182.x
20200510Dichromated gel holograms, 183.x
20200518Dichromated gel holograms, 184.x
20200523dichromated gel holorams 185.x
20200524Squirls on the post, and dichromated gel holograms 186.x.
20200525dichromated gel holograms 187.x
20200530dichromated gel holograms 188.x
20200606Dichromated gel holograms 185.2R2 188.1R1, fox in the back yard, 189.x
20200607Dichromated gel holograms 190.x
20200613Dichromated gel holograms 191.x
20200614Dichromated gel holograms, 192.x
20200628Dichromated gel holograms, 193.x
20200704Thermal images of LED cube, plate coating, laser cooling
20200814Failed attempt at making silver halide film. Boiler lurking in garage.
20200829Alignment images, making holograms
20200906Mirror mounts, and a bunch of test images debugging an extremely sketchy flash adapter.
20200907Pictures and video of holograms 200.1 to 200.3.
20200913Assembly of high voltage flash adapter, test shots
20200919Aligning the analyzing etalon with the lumonics laser, and generally annoying myself. Replaces axial mode aperature,to reduce axial mode issues. Still ring shaped...
20200920Bunch of alignment images for the pulsed apparatus, then one of the pockel cell.
2020092001Lumonics laser fun with the pockel cell alignment.
20200927Arrivial of new washing machine, attempting to use analysing etalon on lumonics laser.
20200927_justinMore attempting to align the lumonics laser.
2020100302Arrival of laser quantum, box dammage.
2020100303Garage sale.
20201004First light from laser quantum laser.
20201009More attempting to align the lumonics laser.
20201010Arrivial of new red CW laser, patched hole in wall, Observing the bifringence in the oscillator ruby, to determine if it is correctly oriented. It is...
20201011Re-Re-Realigning the lumonics laser. This time with an analysing etalon.
2020101101Aligning the pulsed apparatus, with new more robust mirror clamps.
2020101102Analyzing etalon apparatus for the lumonics laser.
2020101103IR images of the new torus 660 laser.
20201013Interior of blue laser power supply.
20201018Replaced switcher power supply in blue laser, RGB laser combining optics, hologram 203.4 first non-bad 3 color hologram.
20201108Attempt to make hologram of head, door now with enhanced safety.
20201115Alignment of pockels cell, holograms 205.x, and 206.x.
20201126Aligning the lumonics laser, fun with an etalon.
20201212More lumonics alignment fun, hologram 207.3
2020121203Location of geophone sensor, thanksgiving pie, filling pockels cell with FC-770, pulsed apparatus surrounded by measurement equipment.
20201213More pulsed apparatus alignment, lighting tests, looking for stray light hitting the plate.
20201219More pulsed apparatus alignment, and lighting tests.
20201220More alignment, and way too many attempts to get a good picture of hologram 208.2.
20201225Hologram 207.x and 208.x setup, pump test.
20201231Holograms 209.x and 210.x.
20210102Bunch of IR images of lasers and pool pump.
20210103Kringle in the time of covid 19, LED cube.
20210110pump project reference images.
20210124Bent pin on fitlet, interior of box with models.
20210203Reference images for deck dimensions.
20210207Pump controller, serial number 1 shakedown testing.
20210228IR images of pump controller project.
20210307pump controler project images.
20210320Deck Project, the beginning of actual work and the pile of bits.
2021032001Location of the geophone for the second run, and ruby optics.
2021032002DCG holograms 211.x
20210321Hologram 211.2, after sealing it.
20210326Deck contstruction, first 4 days.
20210327Deck construction, day 5, deck is mostly done.
20210410Finishing the deck, starting on the patio, Johnny's light.
20210411Pictures of pile of ISA cards for craigslist.
20210421replacement of the beam splitter optics on the pulsed apparatus, then lots of alignment.
20210422lumonics pulsed apparatus alignment, exposures of holograms 212.1 and 212.2.
2021042201Even more lumonics alignment, and lots of obsessing about lighting.
2021042202Patio construction, the first day, lots of gravel.
2021042203Patio construction day 2, lots of concrete.
20210423More patio construction, the molds have been removed, and the surface sealed. It still needs fill dirt around it.
2021042301Holograms 214.x , ish. Bunch of alignment images, and then unusable pictures of 214.1 and 214.3 reconstructed with 532nm.
20210424hologram 215.2 the master moved... and contrast still sucks.
20210501more ruby apparatus alignment.
20210509Final walkthrough details for the deck and patio, pile of tapes, and old JK controls.
20210523First attempt at making cyanotypes.
20210601Pulsed apparatus alignment, and trying different focal length lenses in the illumination beams.
20210705Pictures of holograms of cat, SCSI controller,
20210710pictures of hologam of Eddybear and Sarah
20210717Making of hologram 220.1, emulsion of holograms 220.3 and 220.4, fallen off in the wash.
20210718Hologram 220.1
20210724More lumonics alignment, making of bigfoot hologram.
20210829Incorrect cable from adorama.
20211002Harddisk shipping and packaging from NewEgg.
20211030Installation of defective seagate disks in popcorn. (Due to shingled magnetic recording, they are more accurately described as fixed tape drives.)
20211105Clock radio, and U4 plates.
20220124Completion of the oscilloscope cart.
20220212RGB lasers mode structure, as shown by an etalon. The etalon is not reflective to 457 nm, so no data for that wavelength.
20220312Attempt to make 8x10 color hologram, camera lenses
20220323pictures of surplus motors, pump, editing bin, tests
20220425coin crusher schematic, air conditioner, circuit breakers.
2022042501Covid 19 sign, CWRU Marathon 47.
20220517National watch and clock museum, Columbia PA.
20220521Playing with a jacob's ladder.
20220522NS. Savannah 2022 Maritime day
20220529System Source museum.
20220626Hologram 227.4 illiminated with LED, and the subject illiminiated with the lasers.
20220711Hologram 228.1 2 beam white light color.
20220731pictures of dead tree and branch before and after removal.
20220814Replacing alignment laser on ruby laser, bunch of alignment images, damaged mirrors, peach bits left in my yard by neighbors.
20220821Alignment of pulsed apparatus, lighting tests.
20220825More alignment an lighting tests of pulsed apparatus, making of holding bolex hologram.
20220830Making of hologram 230.1, pulsed H1 of Dan.
20220916Replacing batteries in big UPS, buying way too much candy, new metering pump.
20221106stripped paint on the porch, interior of teryx, mellon found in yard, after painting porch, kitchen waste in neighbor's yard.
20221111Raw plots from joulemeter normal to reference beam, lumonics tests.
2022111101lumonics pulsed apparatus lighting tests.
20221225Pie, videocards, milling machine innards, stamp
20221226Square wave from spindle tach.
20230104fabriacation of wedge clamp, neighbor's fallen tree, first loaf of bread.
20230108fabrication of second LED control box.
2023010801banana experiments
20230117The innards of my heatpump. ABC board.
20230124MFM 2022, 48'th annual CWRU movie marathon
20230128Lumonics pulsed apparatus alignment tests.
20230205Making of holograms 232.1, 232.2, and 233.1
202304302023 physics cookout, and pictures of pool remote display LCD.
2023043001Test images holding camera over microscope camera port.
20230506Couple of microscope test images, 2023 kinetic sculpture race.
20230515Microscope camera test images.
20230521NS Savannah, 2023
20230618holograms 234.1 Dana and Sean, 236.3 Brook and David, 237.2 Ian's Mom.
20230729prototype enclosure for ESP32 module.
20231203System Source museum.
20231231Kringle, cryogenic 316 stainless steel with magnet (did not stick), Squid on mantle.
20240113First successful liquifying of... something. probably has too much O2.
20240115Details of heat pump pressure gauge.
20240122CWRU 49'th movie marathon.
lumonics_images_20171021Images from attempting to align the JK laser.
lumonics_images_20171028More work aligning the lumonics laser.
Valid HTML 4.0!